Nichrome ribbon resistor elements each has a resistance of 1Ω. The element is made from sheet of nichrome alloy, 0.025 cm thick. If the width of the ribbon is 0.3 cm, what length is required per element? Assume specific resistance of nichrome alloy to be 109 micro Ω-cm. (include step-by-step procedures)


1
Expert's answer
2021-10-04T04:55:54-0400

Thicknes a = 0.025 cm

Width of the ribbon b = 0.3 cm

Specific resistance of nichrome alloy = 109 µΩ-cm

Resistance of each element R=1 Ω

Resistance

"R = \frac{ρl}{A} \\

A = 0.025 \times 0.3 = 0.0075 \;cm^2 \\

l = \frac{RA}{ρ} \\

l = \frac{1 \times 0.0075}{109 \times 10^{-6}} = 68.8 \;cm"

Answer: 0.688 m